OpenStudy (anonymous):

Which one of these equations is correctly written in Standard Form? x + y = −4 3x + 2.5y = −4 x + 3y = −12 -3x + y = −12

OpenStudy (dan815):

whats standard form

OpenStudy (anonymous):

i thought it was 

OpenStudy (anonymous):

so all of themm

OpenStudy (anonymous):

but i have to pick one of them

OpenStudy (anonymous):

unless u have a pic for us

OpenStudy (dan815):

lol

OpenStudy (dan815):

ummmmmm i pick a

OpenStudy (anonymous):

do u have the a graph or some sort of extra info?

OpenStudy (anonymous):

those are a b c and d

OpenStudy (dan815):

pick A just do it.. does it tell you the right answer if ur wrong

OpenStudy (anonymous):

Standard Form Ax + By = C

OpenStudy (dan815):

forgive my grammar if u r *

OpenStudy (anonymous):

A. x + y = −4 B. 3x + 2.5y = −4 C. x + 3y = −12 D. -3x + y = −12

OpenStudy (anonymous):

is that all the info u were given, no graph or other equations?

OpenStudy (anonymous):

yes it tells me if im right or wrong but i have 10 questions after this one

OpenStudy (anonymous):

Seems vague.

OpenStudy (anonymous):

if i were u i'd just say my comp froze

OpenStudy (dan815):

pick A and say ur dog did it

OpenStudy (anonymous):

I would go with C

OpenStudy (anonymous):

so if the dog is right, would that mean it is smarter than her?

OpenStudy (anonymous):

lol its online so i cant just say that hahaha thanks though guys!

OpenStudy (anonymous):

ill pick A ;) haha

OpenStudy (anonymous):

Or B

OpenStudy (anonymous):

or Z?

OpenStudy (anonymous):

btw would -3 be less then 1/2?

OpenStudy (dan815):

yes

OpenStudy (dan815):

>_>

OpenStudy (anonymous):

lol yes

OpenStudy (anonymous):

Yeah. I would go with B.

OpenStudy (anonymous):

lol just making sure XD

OpenStudy (dan815):

oh my goodness

OpenStudy (anonymous):

ok ill go with B

OpenStudy (anonymous):

nah A

OpenStudy (anonymous):

Or A